OpenStudy (anonymous):

Given the relationship x2+3y2=12 and dx/dt=2 units/min, find the value of dy/dt at the instant y=1 unit

OpenStudy (anonymous):

so basically your solving a derivative problem right

OpenStudy (anonymous):

Implicit differentiation

OpenStudy (anonymous):

yea same thing

OpenStudy (anonymous):

well the first thing to do would be to take the derivative

OpenStudy (anonymous):

Which would be 2xdx/dt+ 6ydy/dt=0

OpenStudy (anonymous):

and then i think you would plug in all the info you have and solve for dy/dt

OpenStudy (anonymous):

\[x^2+3y^2=12\\2xx'+6yy'=0\] put \(x'=2,y=1\)

OpenStudy (anonymous):

So dy/dt=2/3x

OpenStudy (anonymous):

you can find x also since \[x^2+3y^2=12\] and \[y=1\] you get \[x^2+3=12\\ x^2=9\\ x=\pm 3\]

OpenStudy (anonymous):

Oh, okay. So the answer would be -2. Thank you

OpenStudy (anonymous):

yw